A missile protection system consist of "n" radar sets operation independently,each with a probability of .9 of detecting a missile entering a zone that is covered by all the units.

(a).If "n"=5 and a missile enters the zone,what is the probability that exactly four sets detect the missile?At least one set?

(b).How large must "n" be if we require that the probability of detecting a missile that enters the zone be .999?


1
Expert's answer
2021-04-28T00:45:16-0400

p=0.9, n=5

a) i)Exactly four sets

P(X=4)

"P(X=x) ={n\choose x} p^xq^{n-x}"

"={5 \choose 4}*0.9^4*0.1^1"

=0.3281

ii) Atleast one set

P(X"\ge" 1)

"=\sum_1^5{5\choose x} *0.9^x*0.1^{5-x}"

=0.99999

b) Atleast one radar set should detect the existence of a missile.

"P(X\ge1) =1-P(X=0)"

"0.999=1-0.1^n"

"0.1^n=0.001"

n=3